OpenStudy (anonymous):

f(x)=3x-7 and g(x)=-2x-6 Find (f o g)(4)

OpenStudy (joannablackwelder):

f o g (x) = 3(2x-6)-7 Can you simplify?

OpenStudy (anonymous):

No, I need a step by step explanation

OpenStudy (joannablackwelder):

So, composite functions are when you plug in a function for all the variables in another function. Do you see how I plugged it in?

OpenStudy (anonymous):

Yes, sort of

OpenStudy (joannablackwelder):

I plugged in g(x) for every x in f(x).

OpenStudy (joannablackwelder):

Then you can either simplify or plug in 4 for x in the new function.

OpenStudy (joannablackwelder):

f o g(4) = 3(2*4-6) - 7

OpenStudy (anonymous):

So that would be f o g(4)=-1 right?

OpenStudy (joannablackwelder):

Yup, good job! :)

OpenStudy (anonymous):

Is that it? So all I need to do is put g(x) in for x in f(x), then substitute 4 for x?

OpenStudy (joannablackwelder):

You got it! Just remember that when there are multiple x's in f(x), to plug in g(x) for all of them.

OpenStudy (anonymous):

Thanks, can you help me with a couple more?

OpenStudy (joannablackwelder):

Sure, but in 2 mins.

OpenStudy (anonymous):

Thank you!! What is the inverse of the given relation? y = 3x + 12

OpenStudy (joannablackwelder):

Ok, inverse means switch x and y, then solve for x again.

OpenStudy (joannablackwelder):

So the first step is x = 3y + 12

OpenStudy (joannablackwelder):

Then solve for y.

OpenStudy (joannablackwelder):

Can you do that part?

OpenStudy (anonymous):

You have to get y by itself, so you have to subtract 3 from both sides, right? So it would be x=y+9

OpenStudy (joannablackwelder):

You do want to get y by itself, but subtracting 3 from both sides wouldn't do that. Subtracting 3 from both sides would make x -3 = 3y + 9

OpenStudy (joannablackwelder):

I would suggest first subtracting 12 from both sides to give x - 12= 3y

OpenStudy (joannablackwelder):

Then dividing both sides by 3 to give x/3 - 4 = y

OpenStudy (anonymous):

But how do I solve the other side?

OpenStudy (joannablackwelder):

You can write it the other way. It's the same thing. Is that what you mean

OpenStudy (joannablackwelder):

?

OpenStudy (joannablackwelder):

Like y = x/3 - 4

OpenStudy (anonymous):

That's what I thought, but I second guessed myself

OpenStudy (joannablackwelder):

Yup :)